for the function y=2x2+3, compute the average rate of change between -2 and x=5
Answer by Alan3354(69443)

for the function y=2x2+3, compute the average rate of change between -2 and x=5
---------------
f(x) = 2x^2 + 3
Find f(-2), sub -2 for x
f(-2) = 2(-2)^2 + 3 = 11
------
f(5) = 53
Change in f(x) = 42
Change in x = 7
Avg change = 42/7 = 6
-----------------
PS Use ^ (Shift 6) for exponents.